OpenStudy (anonymous):

Solve using elimination: 5x − 4y − 10 = 0 -3x − 4y = -6 A. (-2, -5) B. (2, 0) C. (-6, -6) D. (6, 5)

OpenStudy (anonymous):

5x - 4y = 10 -3x - 4y = -6 Subtract both the equations: 8x = 16 x = 2 and y = 0.. B is the answer...